Oreo Inc. is a cookie manufacturer making delicious chocolate cookies. Early in 2019, in an effort to produce healthier produWhat is the 2019 Cost of Goods Manufactured for Spinach Oreos

What is the 2019 Cost of Goods Sold for Spinach Oreos

What is the Income from Operations for the Spinach Oreo division

Based on the Income from Operations, was Spinach Oreos a good idea (yes or no)
Answer #1

ANSWER:

(1) Cost of Goods Manufactured = Direct Materials + Direct labor cost + Manufacturing overhead+Beginning WIP Inventory - Ending WIP Inventory

Direct Materials = (5*7,50,000)+ (3*7,50,000) + (0.30* 7,50,000) = 37,50,000 + 22,50,000 + 2,25,000

= 62,25,000

Hours Machine took to produce 7,50,000 units = 7,50,000 / 200 = 3,750

Direct Labor Cost = 0.45 * 7,50,000 = 3,37,500

Factory Overhead cost = 1,250 * No. of Machine Hours = 1,250 * 3,750 = 46,87,500

Ending Inventory WIP = (15,000*5)+(3*15,000)+(0.30*15,000)

= 1,24,500

Cost of Goods Manufactured = 62,25,000 + 3,37,500 + 46,87,500 - 1,24,500

= 1,11,25,500.

(2)Cost Of Goods Sold = Cost of Goods Manufactured - Ending Inventory of Finished goods

Ending Inventory of finished goods = 20,000(7,50,000 - 7,30,000) * 25 = 5,00,000

Cost of Goods Sold = 1,11,25,500 - 5,00,000 = 1,06,25,500

(3)Income From Operations = Gross Profit - Operating Expenses ( Non Manufacturing costs)

Sales - Cost of Goods Sold = Gross Profit = 1,82,50,000 - 1,06,25,500 = 76,24,500

Operating Expenses = Advertising + Promotion + Selling Expenses

= 10,00,000 + 9,00,000 + 56,25,000 = 75,25,000

Income from Operations = 76,24,500 - 75,25,000 = 99,500

(4)Was Spinach Oreos a good Idea: Yes, Spinach Oreo is a good idea since it is landing operational profits to the organisation, and these profits and before taxes henceforth is a great measure for investors to highlight the growth potential of the organizastion. Hence, it was a great idea to manufacture spinach cookies.

Comment for any doubts.